if X and Y are random variables such that Var(X)<and Var(Y)<,show that

Var(aX+bY)=a2Var(X)+b2Var(Y)+2ab cov(X,Y). Where a and b are real constants.


1
Expert's answer
2021-10-27T14:12:48-0400

"E{[(aX+by)-(a\u03bc_X+b\u03bc_Y )]^2 }=E{[a(X-\u03bc_X )+b(Y-\u03bc_Y )]^2 }\n\\\\=E{[a(X-\u03bc_X )]^2 }+E{[2ab(X-\u03bc_X )(Y-\u03bc_Y )]}+E{[b(Y-\u03bc_Y )]^2 }\n\\\\=a^2 \\text{Var\u2061(X)}+2ab\\text{Cov\u2061(X,Y)}+b^2 \\text{Var\u2061(Y)}\n\\\\=a^2 \\text{Var\u2061(X)}+b^2 \\text{Var\u2061(Y)}+2ab\\text{Cov\u2061(X,Y)}"